OpenStudy (anonymous):

x intercept 2 y intercept 2/3 slope intercept form

OpenStudy (anonymous):

well we know that the x intercept is when it crosses the x-axis so y is zero so we have the point (2,0) and y intercept crosses the y-axis so x is zero (0,2/3)

OpenStudy (anonymous):

we know slope intercept form is y=mx+b and b is the y intercept so we can plug in the given there so we have y=mx+2/3

OpenStudy (anonymous):

so we are looking for m or the slope and we know slope is change in y over change in x so we can use the formula\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] for the two points we found out we had and then calculate the slope

OpenStudy (anonymous):

solpe itercept line is what is it 1/6x +2/3

OpenStudy (anonymous):

slope intercept form for the line is

OpenStudy (anonymous):

so it is y2-y1/x2-x1 so that is \[\frac{ \frac{ 2 }{ 3 }-0 }{ 0-2 }\]

OpenStudy (anonymous):

and solve that and you have your slope

OpenStudy (anonymous):

2x+2/3